# Inertial Frame Dragging as a Probe to Differentiate Kerr–Newman Naked Singularities from Black Holes

### 28 Feb 2026 v3

@@ Abstract
We investigate inertial frame dragging and relativistic precession in the Kerr–Newman spacetime and show how gyroscopic observables can operationally discriminate between Kerr–Newman black holes and Kerr–Newman naked singularities. We study a test gyroscope attached to a stationary observer and derive closed-form expressions for the Lense-Thirring, geodetic, and general spin-precession frequencies. A sharp qualitative distinction emerges: for Kerr-Newman black holes, the spin-precession frequency generically diverges as the horizon is approached (remaining finite only for the ZAMO family), whereas for Kerr-Newman naked singularities, the precession remains finite throughout the spacetime, with divergences confined to the ring singularity on the equatorial plane. Working with physically admissible stationary observers (including ZAMOs), we first construct the timelike geodesic motion and the fundamental orbital frequencies for equatorial circular orbits. Using these, we analyse the radial and vertical epicyclic frequencies, the ISCO shift induced by the charge parameter $Q$, and the associated periastron and nodal precession frequencies relevant to quasi-periodic oscillations (QPOs). We demonstrate that nonzero $Q$ produces systematic, and in rapidly rotating regimes nontrivial, modifications to the frequency hierarchy: $\nu_{\rm nod}$ can develop a finite maximum at $r=r_{p}=\mathcal{O}(M)$, its peak amplitude decreases with increasing $Q$, and sign reversals may occur for sufficiently large charge and high spin, signalling a reversal of nodal-precession orientation. These results establish spin-precession behaviour as a robust strong-field probe of horizons versus exposed singularities, with potential implications for testing cosmic censorship using future high-precision precession/QPO measurements.
@@ 2 The spacetime structure
The Kerr-Newman solution in general involves the mass of the black hole $M$, the rotation parameter $a$, the electric charge $Q$ and the spacetime metric (Adamo and Newman,, 2016; Newman et al.,, 1965) is given by
@@
where the quantities $\Delta$ and $\rho^{2}$ have the following explicit expressions,
$\Delta=r^{2}-2Mr+a^{2}+Q^{2};\qquad\rho^{2}=r^{2}+a^{2}\cos^{2}\theta.$
Figure 1: Illustration showing the spin-charge parameter plane of a rotating Kerr-Newman spacetime. In Fig. (a), the curve separates the black hole region from the naked singularity region, corresponding to configurations without an event horizon. Fig. (b) represents the three-dimensional visualisation of the black hole with inner and outer event horizons.
The metric Eq. (1), which extends the Kerr solution by incorporating electric charge, exhibits distinct behaviours based on the values of its parameters, i.e. mass $M$, angular momentum $a$, and electric charge $Q$. The condition $\Delta=0$ determines the locations of the horizons in the Kerr-Newman black hole. Specifically, it defines two distinct three-dimensional hypersurfaces of constant radial coordinate $r$, corresponding to the roots of the equation $\Delta=0$. These roots are given by $r=r_{\pm}=M\pm\sqrt{M^{2}-a^{2}-Q^{2}}$, where $r_{+}$ represents the outer (event) horizon and $r_{-}$ represents the inner (Cauchy) horizon, respectively. The outer horizon, $r_{+}$, is referred to as the event horizon, which marks the boundary beyond which events cannot influence an external observer. The region defined by $r<r_{+}$ is commonly referred to as the interior of the black hole, encompassing both the inner horizon and the central singularity. Specifically, the metric is referred to as ”extremal” when the condition $M^{2}-a^{2}-Q^{2}=0$ is satisfied (Adamo and Newman,, 2016). In this extremal case, the inner and outer event horizons coincide at a single point, located at the radial coordinate $r=M$. However, when $M^{2}-a^{2}-Q^{2}<0$, a different situation arises. In this case, there is no real solution to the equation $\Delta=0$, which determines the locations of the event horizons. Consequently, the space-time does not possess an event horizon to shield the curvature singularity. Without an event horizon, the singularity remains exposed, leading to the formation of what is known as a naked singularity. Therefore, when $M^{2}-a^{2}-Q^{2}<0$, the space-time structure suggests the existence of a naked singularity, a feature of particular relevance to the discussion presented in this paper.
@@
3 Generalised Spin Precession of a Test Gyroscope Around Kerr–Newman Spacetime
@@ 3.1 Behaviour of $\vec{\Omega}_{s}$ at $r=0$
Within this interval, the angular velocity $\Omega$ corresponds to timelike trajectories., Values of $\Omega$ outside this range correspond to spacelike motion and are therefore unphysical. For a static observer outside the ergosphere, we have to put $\Omega=0$; therefore, we get from Eq. (23)
@@ 4.1 Acceleration scalar for stationary observers in Kerr–Newman Spacetime
$\alpha$ $=\sqrt{g^{\beta\gamma}\alpha_{\beta}\alpha_{\gamma}}$ (47)
$=\frac{1}{\rho^{5}\,|K^{2}|}\left\{\Delta\left[M\,(2r^{2}-\rho^{2})\,\mathcal{B}^{2}-r\,\rho^{4}\,\Omega^{2}\sin^{2}\theta\right]^{2}\right.\left.+\,\sin^{2}\theta\,\cos^{2}\theta\,\left[2Mr\,\mathcal{C}^{2}+\Omega^{2}\rho^{4}\Delta\right]^{2}\right\}^{1/2},$
@@
$\alpha=\frac{(r^{2}+a^{2})^{2}-a^{2}\Delta\sin^{2}\theta}{4\,\delta(1-\delta)\,\rho^{7}\,\Delta}\left\{\Delta\left[M(2r^{2}-\rho^{2})\,\mathcal{B}^{2}-r\,\rho^{4}\Omega^{2}\sin^{2}\theta\right]^{2}\right.\\[10.0pt] \left.+\sin^{2}\theta\,\cos^{2}\theta\,\left[2Mr\,\mathcal{C}^{2}+\Omega^{2}\rho^{4}\Delta\right]^{2}\right\}^{1/2}.$ (49)
The above equation specifies the acceleration required to transport a test gyro in the Kerr-Newman spacetime. It reduces to the standerd Kerr result for $Q=0$ as reported in (Chakraborty et al., 2017b, ). This acceleration becomes unbounded, diverging as the event horizon is approached from any direction. By contrast, in the naked singularity case, where no horizon exists, the gyro’s acceleration remains finite throughout the spacetime, including at ($r=0$) for ($\theta\neq\pi/2$), as inferred from the above result.
5 Role of QPOs in Shaping Accretion Flows in Kerr–Newman Geometry
@@
One can readily verify that, in the limiting case $Q\rightarrow 0$, the characteristic frequencies eqs. 65, 66 and 70 smoothly reduce to their standard Kerr counterparts (Chakraborty et al., 2017b, ; Motta et al.,, 2014; Bambi,, 2015). From Figures (9 - 11), it is apparent that, in the black hole regime, the NPPF $\Omega_{nod}$ exhibits a strictly monotonic decay with increasing orbital radius $r$. This behaviour can be summarised by the following condition:
@@
Interestingly, from the $\Omega_{\rm per}(r)$ profiles (see Figure 15), the charge parameter ($Q$) chiefly reshapes the inner-disk behaviour of the accretion disk, and it does so differently for black holes and naked singularities. In the case of black hole, $(\Omega_{\rm per})$ is only relevant outside the black hole ISCO (marked by the dashed line), and it decreases monotonically with radius; increasing ($Q$) shifts the inner edge inward (smaller $(r_{\rm ISCO})$) and correspondingly raises $(\Omega_{\rm per})$ near the disk’s inner rim, but without introducing any qualitative feature such as a turning point or sign change. In contrast, for the naked singularity, the accretion disk can extend to much smaller radii (ISCO of the naked singularity lies much closer in), so $(\Omega_{\rm per})$ reaches significantly larger strong-field values. Moreover, increasing $(Q)$ produces a pronounced inner-region modification, i.e., the naked singularity curve develops a non-monotonic segment and can even exhibit a sign reversal (a narrow region of $(\Omega_{\rm per}<0)$) close to the inner edge, indicating that the periastron-precession sense can flip in the deepest part of the naked singularity disk. Across the panels from Figure (16), the vertical epicyclic (polar) frequency $\Omega_{\theta}(r)$ exhibits a qualitative behaviour that cleanly separates the two geometries. For the black hole configurations (orange curves), ($\Omega_{\theta}$) remains monotonic over the disk domain shown, i.e., it decreases smoothly as ($r$) increases and shows no turning points outside the black hole’s ISCO. In contrast, the naked singularity configurations (blue dashed curves) display a non-monotonic ”dip–rise–decay” profile at low $Q$ value, i.e., a local minimum appears at radii larger than the naked singularity’s ISCO, followed by a mild recovery to a local maximum before joining the common outer fall-off. This ”external minimum” is therefore an intrinsic strong-field signature accessible to the NS disk but absent in the black hole case. Nonetheless, the charge parameter ($Q$) systematically weakens this naked singularity-specific minimum in the displayed profiles. As ($Q$) increases from ($0.1$) to ($0.2$), the dip becomes shallower, and by $Q\simeq(0.4)–(0.6)$ the non-monotonic feature is largely washed out, leaving an almost monotonic naked singularity curve. From the accretion-disk perspective, this means that at small $(Q)$ the naked singularity disk (whose inner edge sits at the ISCO) can extend into a radial band where ($\Omega_{\theta}$) develops the minimum, imprinting a potentially observable spectral/timing feature. Yet at larger ($Q$), that diagnostic becomes progressively less pronounced, since the ($\Omega_{\theta}$) profile smooths out even though the disk still reaches deeper radii than in the black hole case.
@@ Acknowledgments
The author AKC thanks ICARD, Gurukula Kangri (Deemed to be University), Haridwar, India, for providing the necessary facilities and support.

### 22 Mar 2026 v5

@@ Abstract
We study here the precession of the spin of a test gyroscope attached to a stationary observer in the Kerr-Newman spacetime, specifically, to distinguish a naked singularity from a black hole. Extending previous work on Kerr, we thoroughly investigate the role of the electric charge $Q$ and its effects on the precession of the spin of a test gyroscope for black hole and naked singularity. For gyros attached to stationary observers with nonzero angular velocity $\Omega$, we derive closed-form expressions for the general spin-precession, Lense-Thirring, and geodetic precession frequencies. For Kerr-Newman black holes, the spin-precession frequency generically diverges as the event horizon is approached from any direction, remaining finite only for the zero-angular-momentum observer (ZAMO) family. In contrast, for Kerr-Newman naked singularities, the precession remains finite throughout the spacetime, with divergences confined solely to the ring singularity on the equatorial plane. We show that $Q$ systematically modifies these behaviours, particularly in rapidly rotating regimes, and that the acceleration scalar of stationary observers further sharpens the black hole/naked singularity distinction. We then investigate the Lense-Thirring (nodal) precession frequency of equatorial circular orbits in accretion disks around Kerr-Newman black holes and naked singularities. For black holes, the nodal frequency decreases monotonically with increasing radius, whereas for naked singularities it increases, attains a finite maximum, and then decreases; for sufficiently large spin and charge it can even change sign, signalling a reversal of the precession direction. We also compute the fundamental orbital frequencies (Keplerian, radial and vertical epicyclic) and the periastron precession frequency, highlighting the role of the electric charge $Q$ on these frequencies, and showing how the charge parameter modifies the innermost stable circular orbit (ISCO) and the characteristic frequency hierarchy. Since these precession frequencies are intimately related to observed quasiperiodic oscillations (QPOs), such features provide a robust strong-field probe to determine whether a given rotating compact object is a black hole or a naked singularity, with implications for testing the cosmic censorship conjecture using future high-precision timing measurements.
@@ 2 The Spacetime Structure
The Kerr-Newman solution in general involves the mass of the BH $M$, the rotation parameter $a$, the electric charge $Q$ and the spacetime metric (Adamo and Newman,, 2016; Newman et al.,, 1965) is given by
@@
where the the metric functions $\Delta$ and $\rho^{2}$ have the following explicit expressions:
$\Delta=r^{2}-2Mr+a^{2}+Q^{2},\qquad\rho^{2}=r^{2}+a^{2}\cos^{2}\theta.$
Figure 1: Illustration showing the spin-charge parameter plane of a rotating Kerr-Newman spacetime with $M=1$. In Fig. (i), the curve separates the black hole region from the naked singularity region, corresponding to configurations without an event horizon. Fig. (ii) represents the three-dimensional visualisation of the black hole with inner and outer event horizons.
The metric (1), which extends the Kerr solution by incorporating electric charge, exhibits distinct behaviours based on the values of its parameters, i.e. mass $M$, spin parameter $a$ and electric charge $Q$. The structure of the Kerr-Newman spacetime depends critically on the roots of the equation $\Delta=0$, which determine the locations of the horizons. Solving this equation yields
$r_{\pm}=M\pm\sqrt{M^{2}-a^{2}-Q^{2}},$ (2)
where $r_{+}$ denotes the outer (event) horizon and $r_{-}$ corresponds to the inner (Cauchy) horizon. The event horizon represents the boundary beyond which causal signals cannot escape to an external observer. In contrast, the inner horizon marks a secondary causal boundary within the BH interior. The region $r<r_{+}$, therefore, constitutes the black hole interior, containing both the Cauchy horizon and the central curvature singularity.
The nature of the spacetime is determined by the quantity $M^{2}-a^{2}-Q^{2}$. When $M^{2}>a^{2}+Q^{2}$, the spacetime possesses two distinct horizons and corresponds to a Kerr-Newman black hole. In the extremal case, defined by the condition $M^{2}=a^{2}+Q^{2}$ (Adamo and Newman,, 2016), the inner and outer horizons coincide at $r=M$. However, when $M^{2}<a^{2}+Q^{2}$, the equation $\Delta=0$ admits no real solutions and therefore no event horizon exists. In this situation, the curvature singularity is not hidden behind a horizon and becomes visible to distant observers, giving rise to a naked singularity. This classification of the spacetime structure can be clearly illustrated in the spin-charge parameter space shown in Figure (1). In panel (i), the curve defined by the extremality condition $a^{2}+Q^{2}=1$ (with $M=1$) separates the parameter space into two distinct regions: the black hole region, where the inequality $a^{2}+Q^{2}\leq 1$ is satisfied, and the naked singularity region, where $a^{2}+Q^{2}>1$. Thus, configurations lying below the curve correspond to Kerr-Newman black holes with well-defined event horizons, whereas those above the curve represent horizonless spacetimes with exposed singularities. Panel (ii) provides a three-dimensional visualisation of the Kerr-Newman geometry, illustrating the presence of both the inner and outer horizons that characterise the black hole configuration. The parameter-space representation shown in Figure (1) therefore plays a crucial role in distinguishing between physically admissible black hole solutions and naked singularity configurations. This distinction is particularly relevant for the analysis presented in this chapter, since the behaviour of frame dragging and gyroscopic precession differs qualitatively depending on whether the spacetime contains an event horizon or corresponds to a naked singularity.
@@
3 Generalised Spin Precession of a Test Gyroscope in Kerr-Newman Spacetime
@@ 3.1 Behaviour of $\vec{\Omega}_{s}$ at $r=0$
For values of $\Omega$ lying within this interval, the corresponding trajectories are timelike, whereas values outside this range lead to spacelike motion and are therefore unphysical. For a static observer outside the ergosphere, we have to put $\Omega=0$; therefore, we get from Eq. (24)
@@ 4.1 Acceleration scalar for stationary observers in Kerr–Newman Spacetime
$\alpha=\sqrt{g^{\beta\gamma}\alpha_{\beta}\alpha_{\gamma}}=\frac{1}{\rho^{5}\,|K^{2}|}\left\{\Delta\left[\bigl(M(2r^{2}-\rho^{2})-rQ^{2}\bigr)\,\mathcal{B}^{2}-r\,\rho^{4}\,\Omega^{2}\sin^{2}\theta\right]^{2}\right.\\[6.0pt] \left.+\,\sin^{2}\theta\,\cos^{2}\theta\,\left[(2Mr-Q^{2})\,\mathcal{C}^{2}+\Omega^{2}\rho^{4}\Delta\right]^{2}\right\}^{1/2},$ (48)
@@
$\alpha=\frac{(r^{2}+a^{2})^{2}-a^{2}\Delta\sin^{2}\theta}{4\,\delta(1-\delta)\,\rho^{7}\,\Delta}\left\{\Delta\left[\bigl(M(2r^{2}-\rho^{2})-rQ^{2}\bigr)\mathcal{B}^{2}-r\,\rho^{4}\Omega^{2}\sin^{2}\theta\right]^{2}\right.\\[8.0pt] \left.+\sin^{2}\theta\,\cos^{2}\theta\,\left[(2Mr-Q^{2})\,\mathcal{C}^{2}+\Omega^{2}\rho^{4}\Delta\right]^{2}\right\}^{1/2}.$ (50)
The above equation specifies the proper acceleration required to transport a test gyroscope along a stationary trajectory in the Kerr-Newman spacetime. In the limit $Q\to 0$, it reduces exactly to the corresponding Kerr expression reported in (Chakraborty et al., 2017b, ). Although the electric charge introduces additional $Q^{2}$-dependent contributions, the qualitative behaviour of the acceleration remains similar to that in the Kerr geometry. For the black hole case, the acceleration diverges as the event horizon is approached from any direction. This reflects the fact that an observer attempting to remain stationary arbitrarily close to the horizon must undergo an increasingly large proper acceleration. Hence, a stationary timelike worldline cannot be maintained on the horizon with finite acceleration. On the other hand, in the naked singularity case, where no horizon exists, the acceleration remains finite throughout the spacetime except at the ring singularity. In particular, it stays finite even at $r=0$ for $\theta\neq\pi/2$. The divergence appears only at the ring singularity $r=0$ and $\theta=\pi/2$, where $\rho^{2}=0$. Thus, the behaviour of the gyroscope’s acceleration clearly distinguishes the Kerr-Newman black hole from the Kerr-Newman naked singularity.
It is also instructive to note that the proper acceleration vanishes when the gyroscope follows a geodesic trajectory. In the equatorial plane $(\theta=\pi/2)$, Eq. (50) reduces to
$\alpha=\frac{(r^{2}+a^{2})^{2}-a^{2}\Delta}{4\delta(1-\delta)r^{7}\Delta}\left\{\Delta\left[r(Mr-Q^{2})(1-a\Omega)^{2}-r^{5}\Omega^{2}\right]^{2}\right\}^{1/2}.$ (51)
Therefore, the condition $\alpha=0$ yields
$(Mr-Q^{2})(1-a\Omega)^{2}-r^{4}\Omega^{2}=0,$ (52)
whose solutions are
$\Omega_{kep}=\Omega_{\rm\phi}=\frac{\pm\sqrt{Mr-Q^{2}}}{r^{2}\pm a\sqrt{Mr-Q^{2}}},$ (53)
which is precisely the Keplerian frequency (see Eq. (69)) for equatorial circular geodesics in the Kerr-Newman spacetime. Hence, when the gyroscope moves along a circular geodesic with angular velocity $\Omega=\Omega_{kep}=\Omega_{\rm\phi}$, no external force is required to maintain its motion, and the proper acceleration vanishes identically. Thus, Eq. (50) not only characterises the acceleration needed to hold a gyroscope on a stationary orbit in the Kerr-Newman spacetime but also consistently reproduces the geodesic limit in which the motion becomes force-free.
5 Role of Quasi-Periodic Oscillations in Shaping Accretion Flows in Kerr–Newman Geometry
@@
One can readily verify that, in the limiting case $Q\rightarrow 0$, the characteristic frequencies eqs. 69, 70 and 74 smoothly reduce to their standard Kerr counterparts (Chakraborty et al., 2017b, ; Motta et al.,, 2014; Bambi,, 2015). From Figures (9 - 11), it is apparent that, in the black hole regime, the NPPF $\Omega_{nod}$ exhibits a strictly monotonic decay with increasing orbital radius $r$. This behaviour can be summarised by the following condition:
@@
Interestingly, from the $\Omega_{\rm per}(r)$ profiles (see Figure 15), the charge parameter ($Q$) chiefly reshapes the inner-disk behaviour of the accretion disk, and it does so differently for black holes and naked singularities. In the case of black hole, $(\Omega_{\rm per})$ is only relevant outside the black hole ISCO (marked by the dashed line), and it decreases monotonically with radius; increasing ($Q$) shifts the inner edge inward (smaller $(r_{\rm ISCO})$) and correspondingly raises $(\Omega_{\rm per})$ near the disk’s inner rim, but without introducing any qualitative feature such as a turning point or sign change. In contrast, for the naked singularity, the accretion disk can extend to much smaller radii (ISCO of the naked singularity lies much closer in), so $(\Omega_{\rm per})$ reaches significantly larger strong-field values. Moreover, increasing $(Q)$ produces a pronounced inner-region modification, i.e., the naked singularity curve develops a non-monotonic segment and can even exhibit a sign reversal (a narrow region of $(\Omega_{\rm per}<0)$) close to the inner edge, indicating that the periastron-precession sense can flip in the deepest part of the naked singularity disk. Across the panels from Figure (16), the vertical epicyclic (polar) frequency $\Omega_{\theta}(r)$ exhibits a qualitative behaviour that cleanly separates the two geometries. For the black hole configurations (orange curves), ($\Omega_{\theta}$) remains monotonic over the disk domain shown, i.e., it decreases smoothly as ($r$) increases and shows no turning points outside the black hole’s ISCO. In contrast, the naked singularity configurations (blue dashed curves) display a non-monotonic ”dip–rise–decay” profile at low $Q$ value, i.e., a local minimum appears at radii larger than the naked singularity’s ISCO, followed by a mild recovery to a local maximum before joining the common outer fall-off. This ”external minimum” is therefore an intrinsic strong-field signature accessible to the NS disk but absent in the black hole case. Nonetheless, the charge parameter ($Q$) systematically weakens this naked singularity-specific minimum in the displayed profiles. As ($Q$) increases from ($0.1$) to ($0.2$), the dip becomes shallower, and by $Q\simeq(0.4)--(0.6)$ the non-monotonic feature is largely washed out, leaving an almost monotonic naked singularity curve. From the accretion-disk perspective, this means that at small $(Q)$ the naked singularity disk (whose inner edge sits at the ISCO) can extend into a radial band where ($\Omega_{\theta}$) develops the minimum, imprinting a potentially observable spectral/timing feature. Yet at larger ($Q$), that diagnostic becomes progressively less pronounced, since the ($\Omega_{\theta}$) profile smooths out even though the disk still reaches deeper radii than in the black hole case.
@@ 6 Discussion & Summary
    The acceleration scalar provides a clear physical characterisation of stationary motion in the Kerr–Newman spacetime. The derived expression extends the Kerr result by incorporating the effect of charge and reduces smoothly to the known Kerr form in the limit $Q\to 0$. Its behaviour sharply distinguishes the black hole and naked singularity cases: the acceleration diverges at the event horizon for a Kerr–Newman black hole, whereas for a Kerr–Newman naked singularity it remains finite everywhere except at the ring singularity. Moreover, in the equatorial plane, the vanishing of the acceleration reproduces the Keplerian angular velocity of neutral circular geodesics, thereby confirming the consistency of the formalism in the geodesic limit. Hence, the acceleration scalar acts as an effective probe of both the causal structure and the dynamical properties of the Kerr–Newman spacetime.
  • (vi)
@@ Acknowledgments
AKC acknowledge the support of ICARD, Gurukula Kangri (Deemed to be University), Haridwar, India, for providing the necessary facilities.